# FlexProofs: A Vector Commitment with Flexible Linear Time for Computing All Proofs

### v1

@@ 3 A Functional Commitment with Batch Opening
Bünz et al. [BMM+21] proposed a known-exponent multi-exponentiation inner product argument (MIPA) which employs structured-key variants of the commitment in [AFG+16] to commit to $({\bf A},{\bf b})$ and allows one to prove that a value $U\in\mathbb{G}_{1}$ is equal to $\langle{\bf A},{\bf b}\rangle$ against this commitment, where ${\bf A}\in\mathbb{G}_{1}^{n}$ is hidden and ${\bf b}\in\mathbb{F}_{p}^{n}$ is publicly known and structured of the form $(1,b,\ldots,b^{n-1})$ for some $b\in\mathbb{F}_{p}$.
@@
Faster verification. The FC scheme uses the same approach as the known-exponent MIPA in Bünz et al. [BMM+21], which allows the verifier to outsource the computation of $\bf v$$_{\ell}$ to the untrusted prover and reduces verification time. However, this comes at the cost of additionally relying on the $n$-SDH assumption (Assumption 1). We implicitly assume the faster verifier.
@@ 4.1 FlexProofs
In the first layer, each $\mathbf{m}_{j}$ ($j\in[0,\mu)$) is encoded using its multilinear extension $f_{j}:\mathbb{F}_{p}^{\log\mu}\to\mathbb{F}_{p}$ and committed via a commitment scheme $\sf PC$ [PST13] for multilinear polynomials. In the second layer, the vector of the commitments of all ${\bf m}_{j}(j\in[0,\mu))$ is committed using the scheme $\sf FC$ from Section 3. Specifically, the algorithms $\sf VC.Setup$} and $\sf VC.Commit$} in our VC scheme are as follows.
@@
Given the commitment process defined by (8) and (9), there is a two-step method of proving the correctness of ${\bf m}[i]$ for each $i\in[0,N)$. First, use our FC scheme to prove the correctness of $C_{\lfloor i/\mu\rfloor}$ with respect to $C$, i.e., $C_{\lfloor i/\mu\rfloor}=\langle{\bf C},{\bf u}_{\lfloor i/\mu\rfloor}\rangle$. Second, use the PC scheme to prove the correctness of ${\bf m}[i]$ with respect to $C_{\lfloor i/\mu\rfloor}$, i.e., $f_{\lfloor i/\mu\rfloor}(i\bmod\mu)={\bf m}[i]$. Since both steps have time complexity ${\cal O}(\sqrt{N})$, the computational cost of the two-step method is ${\cal O}(\sqrt{N})$. Consequently, proving the correctness of all elements of $\bf m$} one after the other with the two-step method requires ${\cal O}(N\sqrt{N})$ time. By examining the two-step proving process, we observe that: (1) proving the correctness of $C_{j}$ with respect to $C$ takes place $\mu$ times and can be done only once for all elements in ${\bf m}_{j}$; (2) for all $j\in[0,\mu)$, the $\mu$ processes of proving the correctness of all elements in ${\bf m}_{j}$ with respect to $C_{j}$ can be combined by invoking $\sf PC.HyperEval$ on $f_{j}$. Based on these observations, we propose an OpenAll protocol, in which the prover can prove the correctness of all elements in $\bf m$} in ${\cal O}(N)$ time.
@@ 4.2 Linear-Time OpenAll
In this section, we first present OpenAll as an interactive protocol and then make it non-interactive with the Fiat-Shamir variant of [ZXH+22], yielding the algorithms $\sf VC.OpenAll$ and $\sf VC.Verify$. In the non-interactive version of our VC scheme, the executions of the algorithms $\sf VC.Commit$ and $\sf VC.OpenAll$ form a pre-processing phase and the latter algorithm generates $N$ opening proofs for any vector $\bf m$ of length $N$. In particular, the $N$ opening proofs can be used to directly reply any query to $\sf VC.Open$, which is an algorithm different from $\sf VC.OpenAll$ and every time generates an opening proof for any single element of the vector.
@@
Intuition of our protocol. Our OpenAll protocol consists of two steps: (1) proving the correctness of $C_{j}$ with respect to $C$ for all $j\in[0,\mu)$, and (2) combining the $\mu$ processes of proving the correctness of all elements in ${\bf m}_{j}$ with respect to $C_{j}$. For the first step, if the prover uses the scheme $\sf FC$ from Section 3 to generate a proof for each $C_{j}$, then it needs to generate $\mu$ proofs and thus performs $\mathcal{O}(N)$ field operations and $\mathcal{O}(N)$ cryptographic operations (e.g., group exponentiations or pairings) according to Table 1. To reduce this cost, the prover partitions the vector $\bf C$} into blocks and uses $\sf FC$ to generate one batch proof per block. With a batch size of $b$, the prover only needs to generate $\lceil\mu/b\rceil$ batch proofs, which require $\mathcal{O}(N)$ field operations and $\mathcal{O}(N/b)$ cryptographic operations. For the second step, we combine the $\mu$ processes by invoking $\sf PC.HyperEval$ to fold $\mu$ polynomials into a single polynomial [GWC19, BDF+21], called folded polynomial, which is essentially a linear combination of the $\mu$ polynomials, such that if an evaluation claim at a hypercube point is wrong in one of the $\mu$ polynomials, then it will be wrong for the folded polynomial with overwhelming probability.
@@
Making our protocol non-interactive. Up to this point, our protocol has been presented in the interactive setting: in Step 1, the prover sends a message to the verifiers; in Step 2, the prover receives their challenges and responds accordingly. To make it non-interactive, we adopt the Fiat-Shamir variant of [ZXH+22], which is designed for the multi-verifier setting and proven secure in ROM. Our protocol becomes non-interactive as follows. For Step 2, the prover builds a Merkle tree by hashing the coefficients of all polynomials to be folded and their corresponding commitments. Each verifier then receives a proof that includes a Merkle tree membership path, showing that its data is correctly included.
The correctness and security of FlexProofs are proven in Appendix 0.C. A zero-knowledge variant of OpenAll can also be obtained using standard techniques [CFS17, L21] (see Appendix 0.A for the formal definition). In this case, the PC scheme is replaced with its zero-knowledge version, whereas the FC scheme does not require zero-knowledge since it only operates on already hidden commitments.
@@ 4.4 Compatible with a family of zkSNARKs
Paper [PPP25] proposes a construction that combines a VC scheme with a zkSNARK in order to prove the correctness of computations over data originating from multiple sources. To adapt FlexProofs to this construction, we first extend our OpenAll for sub-arrays and then show how FlexProofs is directly compatible with a family of zkSNARKs.
@@
Compatibility with a family of zkSNARKs. According to [PPP25], if a VC scheme encodes the committed vector $\bf m$} as a multi-linear polynomial and supports generation of an evaluation proof for the polynomial at a random point, it is directly compatible with a family of zkSNARKs that encode inputs as multi-linear polynomials [XZZ+19, S20, CBB+23]. Since FlexProofs is built on PC and FC schemes, it can be viewed as encoding the committed vector $\bf m$} into a multi-linear polynomial and supports generating an evaluation proof for the committed polynomial at a random point. We next show how FlexProofs produces such a proof. By Eq. (1), the multi-linear extension of a vector ${\bf m}\in\mathbb{F}_{p}^{N}$ is
@@
Therefore, to prove the evaluation of $f({\bf x})$ at a random point $\bf r$}, the prover sends $C_{F}$ along with proofs generated by the FC and PC schemes to the verifier, who then checks both proofs. The above process can be adapted into a zero-knowledge version via standard techniques [CFS17, L21]. Here, the underlying PC is replaced with its zero-knowledge variant, while the FC component does not require zero-knowledge, as it only operates on hidden commitments.
@@ 5.2 Evaluation of FlexProofs and comparison with HydraProofs
We microbenchmark FlexProofs with batch sizes $b\in\{2\log N,\log^{2}N\}$. We also compared it with HydraProofs [PPP25], the state-of-the-art vector commitment scheme. HydraProofs is the only known VC construction that achieves ${\cal O}(N)$ time to generate all opening proofs for a vector of size $N$, while also being directly compatible with a family of zkSNARKs. For fairness, we implement both schemes in C++ using the same elliptic curve and techniques for $N\in\{2^{16},2^{18},2^{20},2^{22},2^{24}\}$. The experimental results are shown in Table 3.
@@ Appendix 0.B Our FC Scheme
Assumption 1 ($q$-Strong Diffie-Hellman assumption ($q$-SDH)[BB08])
@@
Clearly, if Eq. (2) holds, then Eq. (3) holds. For the other direction, assume the correct value of ${\bf A}^{{\bf b}^{(i)}}$ is $\hat{y}_{i}$ for $i\in[0,t)$, then there exists at least one $j$ such that $y_{j}\neq\hat{y}_{j}$ and $\prod_{i\in[0,t)}y_{i}^{r_{i}}=\prod_{i\in[0,t)}\hat{y}_{i}^{r_{i}}$. Define the $a_{i}=\log_{g_{1}}y_{i}/\hat{y}_{i}$, then there will be at least one $j$ for which $a_{j}\neq 0$ and $\sum_{i\in[0,t)}a_{i}{r_{i}}=0\mod p.$ By the Schwartz-Zippel lemma [S80, Z79], this occurs with probability $\leq 1/p$. Therefore, if Eq. (2) is not fully satisfied, Eq. (3) holds with probability at most 1/p1/p.∎
Non-interactive argument of knowledge in the ROM. Bünz et al. [BMM+21] define a non-interactive argument of knowledge in the random oracle model (ROM). In ROM, a hash function is replaced by a random function which is sampled from the space of all random functions $\rho\leftarrow\mathcal{U}(1^{\lambda})$. The non-interactive argument is an argument system where the prover sends a single message $\pi$, and the verifier using the proof accepts or rejects. Both the prover and verifier have access to a random oracle $\rho$. An argument of knowledge in the ROM has the property that for each convincing prover there exists an extractor which can rewind the prover and reinitialize the random oracle with new randomness.
@@
(Non-interactive argument of knowledge in the ROM, from [BMM+21]). A non-interactive argument is an argument of knowledge for a relation $\mathcal{R}$ with knowledge error $\kappa(\lambda)$ if for every adversary $\widetilde{\mathcal{P}}$ there exists an extractor $\mathcal{E}$ such that
@@ 0.B.1 Function Binding
The existence of $\mathcal{E}_{S}$ is guaranteed by [BMM+21].
@@ Appendix 0.C Correctness and Security Analysis of FlexProofs
Thus, the verifier accepts the proof with probability 1, confirming the correctness of the interactive protocol. And under the random oracle model, the correctness of the non-interactive protocol resulting from the Fiat-Shamir transformation [ZXH+22] follows directly from the established correctness of the underlying interactive scheme.
@@
$\epsilon_{2}$ is negligible. The soundness of the folding scheme for $\sf PC.HyperEval$} is guaranteed by [GWC19, BDF+21]. If $\epsilon_{2}$ is non-negligible, we construct an adversary $\mathcal{B}_{FH}$ that breaks the soundness of the folding scheme with probability at least $\epsilon_{2}$ and thus give a contradiction.
@@
We have shown that the interactive version of FlexProofs satisfies soundness, captured by the position binding property. Next, we show that the interactive version of FlexProofs satisfies a stronger notion of soundness called state restoration soundness [BCS16]. To prove state restoration soundness, we consider adversaries that are allowed to reset or rewind the verifier, potentially reusing the verifier’s randomness or internal state. We note that the above soundness proof is robust to such behaviors. In particular, the reductions used in both cases (function binding and folding soundness) remain valid even if the verifier’s challenge randomness is reused across different transcripts. Hence, the soundness proof already implies that the interactive protocol satisfies state restoration soundness.
The work [ZXH+22] formally proves that as long as an interactive proof protocol is secure against state restoration attacks, then the non-interactive protocol by applying their Fiat-Shamir transformation is sound in the random oracle model. Therefore, we conclude that our non-interactive scheme satisfies soundness (i.e., position binding) in the random oracle model. ∎

### v2

@@ 3 A Functional Commitment with Batch Opening
Bünz et al. [12] proposed a known-exponent multi-exponentiation inner product argument (MIPA) which employs structured-key variants of the commitment in [2] to commit to $({\bf A},{\bf b})$ and allows one to prove that a value $U\in\mathbb{G}_{1}$ is equal to $\langle{\bf A},{\bf b}\rangle$ against this commitment, where ${\bf A}\in\mathbb{G}_{1}^{n}$ is hidden and ${\bf b}\in\mathbb{F}_{p}^{n}$ is publicly known and structured of the form $(1,b,\ldots,b^{n-1})$ for some $b\in\mathbb{F}_{p}$.
@@
Faster verification. The FC scheme uses the same approach as the known-exponent MIPA in Bünz et al. [12], which allows the verifier to outsource the computation of $\bf v$$_{\ell}$ to the untrusted prover and reduces verification time. However, this comes at the cost of additionally relying on the $n$-SDH assumption (Assumption 1). We implicitly assume the faster verifier.
@@ 4.1 FlexProofs
In the first layer, each $\mathbf{m}_{j}$ ($j\in[0,\mu)$) is encoded using its multilinear extension $f_{j}:\mathbb{F}_{p}^{\log\mu}\to\mathbb{F}_{p}$ and committed via a commitment scheme $\sf PC$ [53] for multilinear polynomials. In the second layer, the vector of the commitments of all ${\bf m}_{j}(j\in[0,\mu))$ is committed using the scheme $\sf FC$ from Section 3. Specifically, the algorithms $\sf VC.Setup$} and $\sf VC.Commit$} in our VC scheme are as follows.
@@
Given the commitment process defined by (8) and (9), there is a two-step method of proving the correctness of ${\bf m}[i]$ for each $i\in[0,N)$. First, use our FC scheme to prove the correctness of $C_{\lfloor i/\mu\rfloor}$ with respect to $C$, i.e., $C_{\lfloor i/\mu\rfloor}=\langle{\bf C},{\bf u}_{\lfloor i/\mu\rfloor}\rangle$. Second, use the PC scheme to prove the correctness of ${\bf m}[i]$ with respect to $C_{\lfloor i/\mu\rfloor}$, i.e., $f_{\lfloor i/\mu\rfloor}({\sf Bin}(i\bmod\mu))={\bf m}[i]$. Since both steps have time complexity ${\cal O}(\sqrt{N})$, the computational cost of the two-step method is ${\cal O}(\sqrt{N})$. Consequently, proving the correctness of all elements of $\bf m$} one after the other with the two-step method requires ${\cal O}(N\sqrt{N})$ time. By examining the two-step proving process, we observe that: (1) proving the correctness of $C_{j}$ with respect to $C$ takes place $\mu$ times and can be done only once for all elements in ${\bf m}_{j}$; (2) for all $j\in[0,\mu)$, the $\mu$ processes of proving the correctness of all elements in ${\bf m}_{j}$ with respect to $C_{j}$ share a similar structure and thus can be combined. Based on these observations, we propose an OpenAll protocol, in which the prover can prove the correctness of all elements in $\bf m$} in ${\cal O}(N)$ time.
@@ 4.2 Linear-Time OpenAll
In this section, we first present OpenAll as an interactive protocol and then make it non-interactive with the Fiat-Shamir variant of [74], yielding the algorithms $\sf VC.OpenAll$ and $\sf VC.Verify$. In the non-interactive version of our VC scheme, the executions of the algorithms $\sf VC.Commit$ and $\sf VC.OpenAll$ form a pre-processing phase and the latter algorithm generates $N$ opening proofs for any vector $\bf m$ of length $N$. In particular, the $N$ opening proofs can be used to directly reply any query to $\sf VC.Open$, which is an algorithm different from $\sf VC.OpenAll$ and every time generates an opening proof for any single element of the vector.
@@
Intuition of our protocol. Our OpenAll protocol consists of two steps: (1) proving the correctness of $C_{j}$ with respect to $C$ for all $j\in[0,\mu)$, and (2) combining the $\mu$ processes of proving the correctness of all elements in ${\bf m}_{j}$ with respect to $C_{j}$. For the first step, if the prover uses the scheme $\sf FC$ from Section 3 to generate a proof for each $C_{j}$, then it needs to generate $\mu$ proofs and thus performs $\mathcal{O}(N)$ field operations and $\mathcal{O}(N)$ cryptographic operations (e.g., group exponentiations or pairings) according to Table 1. To reduce this cost, the prover partitions the vector $\bf C$} into blocks and uses $\sf FC$ to generate one batch proof per block. With a batch size of $b$, the prover only needs to generate $\lceil\mu/b\rceil$ batch proofs, which require $\mathcal{O}(N)$ field operations and $\mathcal{O}(N/b)$ cryptographic operations. For the second step, we combine the $\mu$ processes by folding $\mu$ polynomials into a single polynomial [27, 10], called folded polynomial, which is essentially a linear combination of the $\mu$ polynomials, such that if an evaluation claim at a hypercube point is wrong in one of the $\mu$ polynomials, then it will be wrong for the folded polynomial with overwhelming probability. For the folded polynomial, we invoke $\sf PC.HyperEval$.
@@
Making our protocol non-interactive. Up to this point, our protocol has been presented in the interactive setting: in Step 1, the prover sends a message to the verifiers; in Step 2, the prover receives their challenges and responds accordingly. To make it non-interactive, we adopt the Fiat-Shamir variant of [74], which is designed for the multi-verifier setting and proven secure in ROM. Our protocol becomes non-interactive as follows. For Step 2, the prover builds a Merkle tree by hashing the coefficients of all polynomials to be folded and their corresponding commitments. Each verifier then receives a proof that includes a Merkle tree membership path, showing that its data is correctly included.
The correctness and security of FlexProofs are proven in Appendix 0.C. A zero-knowledge variant of OpenAll can also be obtained using standard techniques [19, 40] (see Appendix 0.A for the formal definition). In this case, the PC scheme is replaced with its zero-knowledge version, whereas the FC scheme does not require zero-knowledge since it only operates on already hidden commitments.
@@ 4.4 Compatible with a family of zkSNARKs
Paper [54] proposes a construction that combines a VC scheme with a zkSNARK in order to prove the correctness of computations over data originating from multiple sources. To adapt FlexProofs to this construction, we first extend our OpenAll for sub-arrays and then show how FlexProofs is directly compatible with a family of zkSNARKs.
@@
Compatibility with a family of zkSNARKs. According to [54], if a VC scheme encodes the committed vector $\bf m$} as a multi-linear polynomial and supports generation of an evaluation proof for the polynomial at a random point, it is directly compatible with a family of zkSNARKs that encode inputs as multi-linear polynomials [71, 59, 18]. Since FlexProofs is built on PC and FC schemes, it can be viewed as encoding the committed vector $\bf m$} into a multi-linear polynomial and supports generating an evaluation proof for the committed polynomial at a random point. We next show how FlexProofs produces such a proof. By Eq. (1), the multi-linear extension of a vector ${\bf m}\in\mathbb{F}_{p}^{N}$ is
@@
Therefore, to prove the evaluation of $f({\bf x})$ at a random point $\bf r$}, the prover sends $C_{F}$ along with proofs generated by the FC and PC schemes to the verifier, who then checks both proofs. The above process can be adapted into a zero-knowledge version via standard techniques [19, 40]. Here, the underlying PC is replaced with its zero-knowledge variant, while the FC component does not require zero-knowledge, as it only operates on hidden commitments.
@@ 5.2 Evaluation of FlexProofs and comparison with HydraProofs
We microbenchmark FlexProofs with batch sizes $b\in\{2\log N,\log^{2}N\}$. We also compared it with HydraProofs [54], the state-of-the-art vector commitment scheme. HydraProofs is the only known VC construction that achieves ${\cal O}(N)$ time to generate all opening proofs for a vector of size $N$, while also being directly compatible with a family of zkSNARKs. For fairness, we implement both schemes in C++ using the same elliptic curve and techniques for $N\in\{2^{16},2^{18},2^{20},2^{22},2^{24}\}$. The experimental results are shown in Table 3.
@@ Appendix 0.B Our FC Scheme
Assumption 1 ($q$-Strong Diffie-Hellman assumption ($q$-SDH)[9])
@@
Clearly, if Eq. (2) holds, then Eq. (3) holds. For the other direction, assume the correct value of ${\bf A}^{{\bf b}^{(i)}}$ is $\hat{y}_{i}$ for $i\in[0,t)$, then there exists at least one $j$ such that $y_{j}\neq\hat{y}_{j}$ and $\prod_{i\in[0,t)}y_{i}^{r_{i}}=\prod_{i\in[0,t)}\hat{y}_{i}^{r_{i}}$. Define the $a_{i}=\log_{g_{1}}y_{i}/\hat{y}_{i}$, then there will be at least one $j$ for which $a_{j}\neq 0$ and $\sum_{i\in[0,t)}a_{i}{r_{i}}=0\mod p.$ By the Schwartz-Zippel lemma [58, 77], this occurs with probability $\leq 1/p$. Therefore, if Eq. (2) is not fully satisfied, Eq. (3) holds with probability at most 1/p1/p.∎
Non-interactive argument of knowledge in the ROM. Bünz et al. [12] define a non-interactive argument of knowledge in the random oracle model (ROM). In ROM, a hash function is replaced by a random function which is sampled from the space of all random functions $\rho\leftarrow\mathcal{U}(1^{\lambda})$. The non-interactive argument is an argument system where the prover sends a single message $\pi$, and the verifier using the proof accepts or rejects. Both the prover and verifier have access to a random oracle $\rho$. An argument of knowledge in the ROM has the property that for each convincing prover there exists an extractor which can rewind the prover and reinitialize the random oracle with new randomness.
@@
(Non-interactive argument of knowledge in the ROM, from [12]). A non-interactive argument is an argument of knowledge for a relation $\mathcal{R}$ with knowledge error $\kappa(\lambda)$ if for every adversary $\widetilde{\mathcal{P}}$ there exists an extractor $\mathcal{E}$ such that
@@ 0.B.1 Function Binding
The existence of $\mathcal{E}_{S}$ is guaranteed by [12].
@@ Appendix 0.C Correctness and Security Analysis of FlexProofs
Thus, the verifier accepts the proof with probability 1, confirming the correctness of the interactive protocol. And under the random oracle model, the correctness of the non-interactive protocol resulting from the Fiat-Shamir transformation [74] follows directly from the established correctness of the underlying interactive scheme.
@@
$\epsilon_{2}$ is negligible. The soundness of the folding scheme for $\sf PC.HyperEval$} is guaranteed by [27, 10]. If $\epsilon_{2}$ is non-negligible, we construct an adversary $\mathcal{B}_{FH}$ that breaks the soundness of the folding scheme with probability at least $\epsilon_{2}$ and thus give a contradiction.
@@
We have shown that the interactive version of FlexProofs satisfies soundness, captured by the position binding property. Next, we show that the interactive version of FlexProofs satisfies a stronger notion of soundness called state restoration soundness [6]. To prove state restoration soundness, we consider adversaries that are allowed to reset or rewind the verifier, potentially reusing the verifier’s randomness or internal state. We note that the above soundness proof is robust to such behaviors. In particular, the reductions used in both cases (function binding and folding soundness) remain valid even if the verifier’s challenge randomness is reused across different transcripts. Hence, the soundness proof already implies that the interactive protocol satisfies state restoration soundness.
The work [74] formally proves that as long as an interactive proof protocol is secure against state restoration attacks, then the non-interactive protocol by applying their Fiat-Shamir transformation is sound in the random oracle model. Therefore, we conclude that our non-interactive scheme satisfies soundness (i.e., position binding) in the random oracle model. ∎